# UDC: A Unified Neural Divide-and-Conquer Framework for Large-Scale Combinatorial Optimization Problems

### v1

@@ Appendix B Problem & Sub-problem Definition
$C(\boldsymbol{\rho}^{j})=\sum_{t=0}^{|\boldsymbol{\rho}^{j}|-1}d_{x_{t},x_{t+1}}+d_{x_{n_{j}},x_{0}}$
The optimization of sub-min-max-mTSP should encounter the lengths of routes in adjunct sub-problems. For adjunct length $l_{b},l_{e}$, the objective function of sub-min-max-mTSP with $m$ route is calculated with adding $l_{b}$ to the first route length $C(\boldsymbol{\rho}^{1})$ and $l_{e}$ to $C(\boldsymbol{\rho}^{m})$. Similar to the normal min-sum objective function, the min-max objective function can also be decomposed to separate objective functions of sub-problems so UDC can be applied to it. To maintain optimality, the testing stage will not update adjacent sub-min-max-mTSP simultaneously.
@@ C.1 Dividng Stgae: Graph Construction
Lighter-weight GNNs require the construction of a graph input (the sparse graph $\mathcal{G}_{D}$). For the CO problems involved in this paper, some CO problems naturally have graph-like features (e.g., the MIS problem) so their sparse graph $\mathcal{G}_{D}=\mathcal{G}$. However, the inputs of most other CO problems are coordinates or matrics instead. So, for the VRPs, we adopt the graph construction method in heatmap-based solvers [31] and GLOP [13], generally using the KNN method to connect edges in the edge set $\mathbb{E}\in\mathcal{G}_{d}$. For CVRP, OVRP, ATSP, OP, PCTSP, SPCTSP, and min-max mTSP (min-max mTSP specifically uses dummy depots for constraints [67]), we directly use a similar method to link edges based on KNN. In addition, for the KP without neighborhood properties, we use $1-\frac{d^{\prime}_{i,j}}{\max_{q\in{1,\ldots,N}}d^{\prime}_{i,q}}$ as the weight of $(i,j)$ for neighborhood selection where $d^{\prime}_{i,j}=\frac{v_{i}+v_{j}}{w_{i}+w_{j}}$ and use the KNN for edge connection as well. In the construction of all the CO problems (excluding MIS), we set neighborhood $K=100$, so the number of edges in the graph is limited to $|\mathbb{E}|=\mathcal{O}(KN)$.
@@ C.2 Dividing Policy: Anisotropic Graph Neural Networks
In modeling the dividing policy, we follow [31] on the choice of neural architectures. The backbone of the graph neural network is an anisotropic GNN with an edge gating mechanism [40]. Let $\boldsymbol{h}_{i}^{\ell}$ and $\boldsymbol{e}_{i,j}^{\ell}$ denote the $d$-dimensional node and edge features at layer $\ell$ associated with node $i$ and edge $(i,j)$, respectively. The features at the next layer are propagated with an anisotropic message-passing scheme:
@@
where $\boldsymbol{U}^{\ell},\boldsymbol{V}^{\ell},\boldsymbol{P}^{\ell},\boldsymbol{Q}^{\ell},\boldsymbol{R}^{\ell}\in\mathbb{R}^{d\times d}$ are the learnable parameters of layer $\ell$, $\alpha$ denotes the activation function (we use $\mathrm{SiLU}$ [68] in this paper), $\mathrm{BN}$ denotes the Batch Normalization operator [69], $\mathcal{A}$ denotes the aggregation function (we use mean pooling in this paper), $\sigma$ is the sigmoid function, $\odot$ is the Hadamard product, and $\mathcal{N}_{i}$ denotes the outlines (neighborhood) of node $i$. We use a 12-layer ($L$=12) AGNN with a width of $d=64$. The total space and time complexity of AGNN is $\mathcal{O}(KNd)$.
Next, based on the Eq. (4), the final layer edge representation $\boldsymbol{e}_{i,j}^{L}$ of AGNN is used to generate the heatmap. Heatmap-based solvers generally use an MLP to process the $\boldsymbol{e}_{i,j}$. To integrate partial solution-related representations in the heatmap, UDC has designed an additional T-revise method, which will re-generate the heatmap by additional $T-1$ times using MLPs based on the existing partial solutions after multiple heatmap decoding steps. This operation increases the time and space complexity to $\mathcal{O}(TKNd)$, but also significantly improves the effect. Enabling the T-revision method, the total initial solution generation strategy is shown below.
$\pi_{d}(\boldsymbol{x}_{0}|\mathcal{G},\Omega,\phi)$ ={$&p(x_{0,1})\prod_{c=0}^{T}\Big{[}p(\mathcal{H}|\mathcal{G},\Omega,\phi,x_{0,1:nt})\prod_{t=2}^{\tau}\frac{\text{exp}(\mathcal{H}_{x_{0,t-1},x_{0,t}})}{\sum_{i=t}^{N}\text{exp}(\mathcal{H}_{x_{0,t-1},x_{0,i}})}\Big{]},&\text{if}\ \boldsymbol{x}_{0}\in\Omega\\ &0,&\text{otherwise}\end{aligned}\right..$ (17)
where $p(x_{0,1})=\frac{1}{N}$ for CO problems without a depot (i.e., TSP, KP, MIS, ATSP). For other CO problems $x_{0,1}=\text{depot}$ and $p(x_{0,1})=1$. In training UDC, we set $T=\lfloor\frac{N}{n}\rfloor$. In testing, we set $T=\lfloor\frac{N}{n}\rfloor$ for 500-node to 4,999-node random instances and set $T=10$ for very large-scale CO instances and CVRPLib-XXL datasets.
@@ C.3 Conquering Stage: Sub-problem Preparation
The method to decompose CO problems to their corresponding sub-CO problems is detailed in B. Generally, for different problems, when decomposing, we need to 1) Mark the already determined solution (starting point in VRP, point linking visited nodes in MIS) 2) Handle the original constraints of the current sub-problem (capacity of CVRP, capacity of KP, etc.). 3) In testing, handle the margin of the constraints in the original CO problem (i.e., KP, OP, PCTSP, SPCTSP) to ensure the accessibility of the optimal solution.
@@ C.4 Conquering Stage: Normalization
As presented in B, TSP, ATSP, CVRP, OVRP, and min-max mTSP conduct normalization on the coordinate. ATSP normalizes the distance matrix by dividing by the maximal number in each sub-ATSP matrix (i.e., $d^{\prime}_{i,j}=\frac{d_{i,j}}{max_{o,q}\in\{1,\ldots,N\}d_{o,q}}$) and other CO problems normalize following the method in GLOP [13] the rotation example making the x-axis longer and dividing it by the existing length of the x-axis makes the maximum length of the x-axis. Let $(x^{\prime}_{i},y^{\prime}_{i})$ denotes the coordinates of the $i$-th node after transformation; $x_{\max}$, $x_{\min}$, $y_{\max}$, and $y_{\min}$ denote the bounds of an sub-VRP $\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ (i.e, $x_{\max}=\max_{i\leq n}x_{i}$, et al.). Then the coordinate transformation of coordinate $(x_{i},y_{i})$ is formulated as
@@
where the scale coefficient $sc=\frac{1}{\max(x_{\max}-x_{\min},y_{\max}-y_{\min})}$. Besides normalizing the coordinates, sub-SPCTSP and sub-PCTSP normalize the numerical constraint by dividing the collected prize of the current sub-solution.
@@ C.5 Conquering Stage: Conquering Policy
This paper adopts 5 solvers as the conquering policy, including AGNN [35] for MIS, ICAM [16] for TSP, CVRP, OVRP, and KP, POMO [5] for PCTSP, SPCTSP, OP, MatNet [70] for ATSP, and DPN [60] for min-max mTSP. For the conquering policy, we employ 12-layer and 256-dimensional AGNN, 6-layer ICAM without AAFT, 6-layer POMO, and 5-layer MatNet, for these experiments respectively. In training, the learning rate of all these solvers is set to $lr=1e-4$.
The baseline in training conquering policies generally follows the design in POMO [5], but the generation process of the $\beta$ baselines on each CO problem differs. Some CO problems have the trait of symmetric so for VRP whose solution is symmetric (i.e., TSP, CVRP, OVRP, OP, PCTSP, SPCTSP), we sample $\frac{\beta}{2}$ solutions on both the starting and ending node. For other CO problems, we sample $\beta$ solutions from the starting point (i.e., the first randomly selected vertex in MIS).
In testing, the conquering policy in the UDC samples only one sub-solution (i.e., $\beta=1$) for asymmetric sub-CO problems and sampling only two sub-solutions (i.e., $\beta=2$) for symmetric ones.
@@ C.6 Total Algorithm
The total algorithm of training UDC is presented in Algorithm 1. The total algorithm of testing UDC is presented in Algorithm 2.
@@
5:  Sample $\alpha$ initial tours: $\{\boldsymbol{x}_{0}^{1},\ldots,\boldsymbol{x}_{0}^{\alpha}\}$ based on $\pi_{c}$ in Eq. (4). (Eq. (17) for T-revisit enabled.) $\beta\leftarrow 2$ for TSP, ATSP, OP, (S)PCTSP.
@@
C.7 Complexity
@@
Using the lightweight GNN for global partition, both the time and space complexities of UDC are reduced compared to both single-stage SL-based sub-path solvers and single-stage RL-based constructive-based solvers. When sampling $\alpha$ $T$-revisit enabled initial solutions, the UDC method that executes $r$ conquering stages has a time complexity of $\mathcal{O}(\alpha TKN)$ and a space complexity of $\mathcal{O}(\alpha KN)$. Both the time and space complexity of a single conquering stage on a single sub-problem is $\mathcal{O}(n^{2})$. There are $\alpha\lfloor\frac{N}{n}\rfloor$ parallel sub-problems for each instance so the complexity is $\alpha\lfloor\frac{N}{n}\rfloor\mathcal{O}(n^{2})$. However, the $\alpha\lfloor\frac{N}{n}\rfloor$ sub-problems can be processed in parallel (as batch size = $\alpha\lfloor\frac{N}{n}\rfloor$). Considering GPU parallel, the time complexity of UDC can be considered as $\mathcal{O}(\alpha TKN+rn^{2})$ and the space complexity is $\mathcal{O}(\alpha KN+\alpha\lfloor\frac{N}{n}\rfloor n^{2})$.
Sub-path-based constructive solver adopts a decoder-only model structure so for a $\tau$-length solution, the time complexity is $\mathcal{O}(\tau N^{2}d)=\mathcal{O}(N^{3})$ (d is the dimension of embeddings) and the space complexity is $\mathcal{O}(N^{2})$. ICAM is a representative RL-based constructive solver, adopting an encoder-decoder-based model structure. The space complexity is $\mathcal{O}(N^{2})$ and the time complexity is $\mathcal{O}(\tau Nd)=\mathcal{O}(N^{2})$ for single-start ICAM and $\mathcal{O}(\tau N^{2}d)=\mathcal{O}(N^{3})$ for N-start (multi-start) ICAM. The comparison of time and space complexity is listed in Table 6, UDC generally adopts the parameter of $T=10$, $\alpha=1$, $K=100$. So on CO problems with $N\geq 1,000$, the proposed UDC-$\boldsymbol{x}_{2}$ demonstrates less complexity compared to $\mathcal{O}(N^{2})$ and exhibits superior efficiency. Large-scale experiments ($N\geq 5,000$) in Table 11 and Appendix D.4 prove this statement.

### v2

@@ B.2 Definition of Problem & Sub-problem
$C(\boldsymbol{\rho}^{j})=\sum_{t=0}^{|\boldsymbol{\rho}^{j}|-1}d_{\rho^{j}_{t},\rho^{j}_{t+1}}+d_{\rho^{j}_{n_{j}},\rho^{j}_{0}}$
Similar to the normal min-sum objective function, the min-max objective function can also be decomposed to independent objective functions of sub-problems so UDC can be evaluated on it. However, as the difference, the optimization of sub-min-max mTSP should encounter the lengths of sub-routes in adjunct sub-problems. For a sub-min-max mTSP with lengths of sub-routes in adjunct sub-problems being $l_{b}$ and $l_{e}$, the objective function of sub-min-max mTSP solution with $m$ sub-routes $\boldsymbol{s}=\{\boldsymbol{\rho}^{\prime,1},\boldsymbol{\rho}^{\prime,2},\ldots,\boldsymbol{\rho}^{\prime,m}\}$ is calculated after adding $l_{b}$ to the first sub-route length $C(\boldsymbol{\rho}^{1})$ and $l_{e}$ to $C(\boldsymbol{\rho}^{m})$ as follows:
minimize $f(\boldsymbol{s})=\max_{i=1}^{m}\{C(\boldsymbol{\rho}^{\prime,1})+l_{b},C(\boldsymbol{\rho}^{\prime,2}),\ldots,C(\boldsymbol{\rho}^{\prime,m})+l_{e}\},$ (16)
Note that only holding two adjunct sub-problems unchanged can make the optimization of the sub-min-max mTSP consistent with the original min-max mTSP. So, to maintain optimality, the testing stage will not update adjacent sub-min-max-mTSP simultaneously.
@@ C.1 Dividng Stgae: Graph Construction
Light-weight GNNs require constructing a graph as input (the sparse graph $\mathcal{G}_{D}$). Some CO problems involved in this paper naturally have graph-like input (e.g., the MIS problem) so their sparse graph $\mathcal{G}_{D}=\mathcal{G}$. However, the inputs of most other CO problems are coordinates or matrics. So, for the VRPs, we adopt the graph construction method in heatmap-based solvers [31] and GLOP [13], generally using the KNN method to connect edges in the edge set $\mathbb{E}\in\mathcal{G}_{d}$. For CVRP, OVRP, ATSP, OP, PCTSP, SPCTSP, and min-max mTSP (min-max mTSP specifically uses dummy depots for constraints [72]), we directly use a similar method to link edges based on KNN. In addition, for the KP without neighborhood properties, we use $1-\frac{d^{\prime}_{i,j}}{\max_{q\in{1,\ldots,N}}d^{\prime}_{i,q}}$ as the weight of $(i,j)$ for neighborhood selection where $d^{\prime}_{i,j}=\frac{v_{i}+v_{j}}{w_{i}+w_{j}}$ and use the KNN for edge connection as well. In the construction of all the involved CO problems (excluding MIS), we set neighborhood size $K=100$, so the number of edges in the graph is limited to $|\mathbb{E}|=\mathcal{O}(KN)$.
@@ C.2 Dividing Policy: Anisotropic Graph Neural Networks
In modeling the dividing policy, we follow previous heatmap-based methods [39] on the choice of neural architectures. The graph neural network of UDC is an anisotropic GNN with an edge gating mechanism [40]. Let $\boldsymbol{h}_{i}^{\ell}$ and $\boldsymbol{e}_{i,j}^{\ell}$ denote the $d$-dimensional node and edge features at layer $\ell$ associated with node $i$ and edge $(i,j)$, respectively. The features at the next layer are propagated with an anisotropic message-passing scheme:
@@
where $\boldsymbol{U}^{\ell},\boldsymbol{V}^{\ell},\boldsymbol{P}^{\ell},\boldsymbol{Q}^{\ell},\boldsymbol{R}^{\ell}\in\mathbb{R}^{d\times d}$ are the learnable parameters of layer $\ell$, $\alpha$ denotes the activation function (we use $\mathrm{SiLU}$ [73] in this paper), $\mathrm{BN}$ denotes the Batch Normalization operator [74], $\mathcal{A}$ denotes the aggregation function (we use mean pooling in this paper), $\sigma$ is the sigmoid function, $\odot$ is the Hadamard product, and $\mathcal{N}_{i}$ denotes the outlines (neighborhood) of node $i$. We use a 12-layer ($L$=12) AGNN with a width of $d=64$. The total space and time complexity of AGNN is $\mathcal{O}(KNd)$. In training all involved CO problems, the learning rate of the dividing policy is set to $lr=1e-4$.
$T$-revisit initial solution decoding. Next, based on the Eq. (4), the final layer edge representation $\boldsymbol{e}_{i,j}^{L}$ of AGNN is used to generate the heatmap. Heatmap-based solvers generally use an MLP to process the $\boldsymbol{e}_{i,j}$. To integrate partial solution-related representations in the heatmap, UDC has designed an additional $T$-revisit method, which will re-generate the heatmap by an additional $T-1$ times using MLPs based on the existing partial solutions after multiple heatmap decoding steps. This operation increases the time and space complexity to $\mathcal{O}(TKNd)$, but significantly improves the solution quality. Enabling the $T$-revisit method, the total initial solution generation strategy is shown below.
$\pi_{d}(\boldsymbol{x}_{0}|\mathcal{G},\Omega,\phi)$ ={$&p(x_{0,1})\prod_{c=0}^{T-1}\Big{[}p(\mathcal{H}|\mathcal{G},\Omega,\phi,x_{0,1:\lfloor\frac{N}{T}\rfloor c+1})\prod_{t=2}^{\tau}\frac{\text{exp}(\mathcal{H}_{x_{0,t-1},x_{0,t}})}{\sum_{i=t}^{N}\text{exp}(\mathcal{H}_{x_{0,t-1},x_{0,i}})}\Big{]},&\text{if}\ \boldsymbol{x}_{0}\in\Omega\\ &0,&\text{otherwise}\end{aligned}\right..$ (19)
It represents decoding solutions from the heatmap $\mathcal{H}\sim p(\mathcal{H}|\mathcal{G},\Omega,\phi,x_{0,1:\lfloor\frac{N}{T}\rfloor c+1})$ and a new heatmap will generates every $\lfloor\frac{N}{T}\rfloor$ steps. For CO problems without a depot (i.e., TSP, KP, MIS, ATSP), $p(x_{0,1})=\frac{1}{N}$ and for other CO problems, $x_{0,1}=\text{depot}$ and $p(x_{0,1})=1$. In training UDC, we set $T=\lfloor\frac{N}{n}\rfloor$. In testing, we generally set $T=\lfloor\frac{N}{n}\rfloor$ as well for 500-node to 4,999-node random instances. For efficiency, we set $T=10$ for very large-scale CO instances and CVRPLib-XXL datasets.
@@ C.3 Conquering Stage: Sub-problem Preparation
The detail of decomposing CO problems to their corresponding sub-CO problems is shown in B. Generally, to generate the constraints of different sub-CO problems at this step, we need to 1) Keep the already determined solutions and likes unchanged (e.g., starting point in VRP, points connected to visited nodes in MIS) 2) Dispatch the original constraints to the current sub-problems (capacity of CVRP, capacity of KP, maximum length of OP, etc.). Usually, this kind of constraint is calculated from nodes in the original sub-solution. 3) In testing, consider the margin of constraints in the original CO problems (i.e., KP, OP, PCTSP, SPCTSP) to ensure the accessibility of optimal solutions.
@@ C.4 Conquering Stage: Normalization
As presented in B, TSP, ATSP, CVRP, OVRP, and min-max mTSP conduct normalization on the inputs of their sub-problems. ATSP normalizes the distance matrix by dividing by the maximum in each sub-ATSP cost matrix (i.e., $d^{\prime}_{i,j}=\frac{d_{i,j}}{max_{o,q}\in\{1,\ldots,N\}d_{o,q}}$) and other sub-CO problems are normalized following the method in GLOP [13]. Let $(x^{\prime}_{i},y^{\prime}_{i})$ denotes the coordinates of the $i$-th node after normalization, $x_{\max}$, $x_{\min}$, $y_{\max}$, and $y_{\min}$ denote the bounds of an sub-VRP instance $\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ (i.e, $x_{\max}=\max_{i\leq n}x_{i}$, et al.). Then, the coordinate transformation of coordinate $(x_{i},y_{i})$ is formulated as follows:
@@
where the scale coefficient $sc=\frac{1}{\max(x_{\max}-x_{\min},y_{\max}-y_{\min})}$. Besides normalizing the coordinates, sub-SPCTSP and sub-PCTSP normalize prizes of nodes by dividing the total collected prize of the current sub-solution.
@@ C.5 Conquering Stage: Conquering Policy
UDC adopts 5 solvers as the conquering policy, including AGNN [35] for MIS, ICAM [16] for TSP, CVRP, OVRP, and KP, POMO [5] for PCTSP, SPCTSP, OP, MatNet [41] for ATSP, and DPN [63] for min-max mTSP. We employ 12-layer and 256-dimensional AGNN, 6-layer ICAM without AAFT, 6-layer POMO, and 5-layer MatNet, for these experiments respectively. In training, the learning rate of all these solvers is set to $lr=1e-4$ as well.
The baseline calculation method in training conquering policies generally follows POMO [5] (sampling $\beta$ sub-solutions for each sub-CO instance), but the generation process of the $\beta$ sub-solutions is different for each CO problem. Some sub-CO problems (e.g., sub-TSP, sub-CVRP, sub-OVRP, sub-OP) have symmetry (i.e., flipping sub-solutions will not change their objective values) so we enable a two sides conquering baseline, sampling $\frac{\beta}{2}$ solutions on both the starting and ending nodes. For other CO problems, we sample $\beta$ solutions from the starting point (e.g., the first randomly selected vertex in sub-MIS and the starting node of sub-ATSP). Please refer to the second to last line in Table 7 for whether the conquering policy enables the two sides conquering baseline in solving each CO problem. In testing, the conquering policy in the UDC samples only one sub-solution (i.e., $\beta=1$) for asymmetric sub-CO problems and sampling only two sub-solutions (i.e., $\beta=2$) for symmetric ones.
Reason for using different solvers for different CO problems
Using the same model for conquering policy is an ideal choice. However, the experiment of UDC involves 10 CO problems, and some CO problems, like min-max mTSP [63] and ATSP [41], require specific model structures to process their inputs and constraints. Consequently, there is no constructive solver that can be used for all 10 involved CO problems. For each CO problem, UDC adopts the best-performance constructive solver for the conquering policy (i.e., summarized in Table 7).
UDC exhibits flexibility in choosing a constructive solver for conquering policy. For well-developed CO problems like TSP and KP, there are multiple constructive solvers available. Section 5.2 and Appendix E.6 present an ablation study on the selection of these constructive solvers for TSP and KP, respectively. Both sets of results indicate that the final performance of UDC is not sensitive to the model selection for the conquering policy. Therefore, when generalizing UDC to new CO problems, it may not be necessary to specifically choose which available constructive solver to model the conquering policy.
@@ C.6 Total Algorithm
The total algorithm of training UDC is presented in Algorithm 1. The total algorithm of testing UDC on instances is presented in Algorithm 2.
@@
5:  Sample $\alpha$ initial tours: $\{\boldsymbol{x}_{0}^{1},\ldots,\boldsymbol{x}_{0}^{\alpha}\}$ based on $\pi_{c}$ in Eq. (4). (Eq. (19) for $T$-revisit enabled.) $\beta\leftarrow 2$ for TSP, ATSP, OP, PCTSP, SPCTSP.
@@
C.7 MDP of the dividing and conquering stages
In addition to the algorithms in Appendix C.6, this section will provide necessary descriptions of the training objective of both policies and the implicit MDP for the two stages.
It should be noted that the training objectives and MDP are not created originally by UDC, and they are also adopted in a series of neural divide-and-conquer methods [13, 20].
Objective functions of optimizing the two networks
. For a CO problem (or a sub-CO problem) with the objective function $f(\cdot)$, the goal of training both networks is to maximize the reward functions of their corresponding MDP. The reward is $-f(\boldsymbol{x}_{2})$ for the dividing policy (AGNN in UDC), $-f(\boldsymbol{s}^{1})$ for the constructive solver (conquering policy) in the Conquer step (i.e., the first conquering stages), and $-f(\boldsymbol{s}^{2})$ in the Reunion step.
MDP for the dividing stage:
The MDP $\mathcal{M}_{d}=\{\mathcal{S}_{d},\mathcal{A}_{d},\boldsymbol{r}_{d},\mathcal{P}_{d}\}$ of the dividing stage can be represented as follows:
State. State $st_{d}\in\mathcal{S}_{d}$ represents the current partial solution. The state in the $t$-th time step is the current partial solution with $t$ nodes $st_{d,t}=\boldsymbol{x}_{0,t}=(x_{1},x_{2},...,x_{t})$. $st_{d,0}$ is empty and $st_{d,T}=\boldsymbol{x}_{0}$.
Action & Transaction. The action is to select a node at time step $t$, i.e., $a_{d,t}=x_{t+1}$. The chosen node needs to ensure that the partial solution $st_{d,t+1}=(x_{1},x_{2},...,x_{t},x_{t+1})$ is valid (i.e., $st_{d,T}\in\Omega$).
Reward. Every single time step has the same reward $r_{d,t}=-f(\boldsymbol{x}_{2})$ which is the objective function value of the final solution $\boldsymbol{x}_{2}$ after the whole DCR process.
Policy. The policy $\pi_{d}$ is shown in Eq. (4) (Eq. (19) for $T$-revisit enable).
MDP for the conquering stage:
The MDP $\mathcal{M}_{c}=\{\mathcal{S}_{c},\mathcal{A}_{c},\boldsymbol{r}_{c},\mathcal{P}_{c}\}$ of any single conquering stage is represented similarly as follows:
State. Each state $st_{c}\in\mathcal{S}_{c}$ represents the current partial solution. The state in the $t$-th time step is the current partial solution with $t$ nodes $st_{c,t}=\boldsymbol{s}_{t}=(s_{1},s_{2},...,s_{t})$. $st_{c,0}$ is empty and $st_{c,T}=\boldsymbol{s}$.
Action & Transaction. The action is to select a node at time step $t$ as well, i.e., $a_{c,t}=s_{t+1}$.
Reward. The reward in each time step becomes the objective value of sub-CO solution $\boldsymbol{s}$, i.e., $r_{c,t}=-f(\boldsymbol{s})$.
Policy. The policy $\pi_{c}$ is shown in Eq. (5).
C.8 Time & Space Complexity
@@
With a lightweight GNN for global partition and constructive solvers for parallel small-scale sub-problems, both the time and space complexities of UDC are reduced compared to single-stage SL-based sub-path solvers and single-stage RL-based constructive-based solvers. When sampling $\alpha$ $T$-revisit enabled initial solutions ($T$=1 for disabling the $T$-revisit), the dividing stage of UDC has a time complexity of $\mathcal{O}(\alpha TKN)$ and a space complexity of $\mathcal{O}(\alpha KN)$.
Both the time and space complexity of a single conquering stage on a single sub-problem is $\mathcal{O}(n^{2})$. There are $\alpha\lfloor\frac{N}{n}\rfloor$ parallel sub-problems for each instance so the complexity should be $\alpha\lfloor\frac{N}{n}\rfloor\mathcal{O}(n^{2})$. However, the $\alpha\lfloor\frac{N}{n}\rfloor$ sub-problems can be processed in parallel (as batch size = $\alpha\lfloor\frac{N}{n}\rfloor$) in GPUs, so the total time complexity of UDC with $r$ conquering stages can be considered as $\mathcal{O}(\alpha TKN+rn^{2})$ and the space complexity is $\mathcal{O}(\alpha KN+\alpha\lfloor\frac{N}{n}\rfloor n^{2})$.
Sub-path-based constructive solver adopts a decoder-only model structure so for a $\tau$-length solution, the time complexity is $\mathcal{O}(\tau N^{2}d)=\mathcal{O}(N^{3})$ ($d$ is the dimension of embeddings, we regard $\tau=\mathcal{O}(n)$) and the space complexity is $\mathcal{O}(N^{2})$. ICAM is a representative RL-based constructive solver, adopting an encoder-decoder-based model structure. The space complexity is $\mathcal{O}(N^{2})$ and the time complexity is $\mathcal{O}(\tau Nd)=\mathcal{O}(N^{2})$ for single-start ICAM and $\mathcal{O}(\tau N^{2}d)=\mathcal{O}(N^{3})$ for $N$-start (multi-start) ICAM. The comparison of time and space complexity is listed in Table 6, UDC generally adopts the parameter of $T=10$, $\alpha=1$, $K=100$ in testing. So on CO problems with $N\geq 1,000$, the proposed UDC-$\boldsymbol{x}_{2}$ demonstrates less inference complexity compared to $\mathcal{O}(N^{2})$ and exhibits superior efficiency. Empirical results especially very-large-scale experiments ($N\geq 5,000$) in Table 11 and Appendix D.5 prove the above statements.